Can someone please help me figure out how to use a ti 84 calculator to do E, F, G, and H. Thanks in advance.
Answer #1

Solution:-

e) z0 = 0.978

P(-z0 < z < 0) = 0.3351

P(z < z0) = 0.50 - 0.3351

P(z < z0) = 0.1649

z0 = 0.978

f) z0 = 2.772

P(-2 < z < z0) = 0.9738

P(-2 < z < z0) = P(z > - 2) - P(z > z0)

0.9738 = 0.977 - P(z > z0)

P(z > z0) = 0.977 - 0.9738

P(z > z0) = 0.0032

z0 = 2.772

g) z0 = 0

P(z > z0) = 0.50

z0 = 0

h) z0 = - 3.012

P(z < z0) = 0.0013

P(z > z0) = 1 - 0.0013

P(z > z0) = 0.9987

z0 = - 3.012